# A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v1

@@ 3.2 Empirical Studies
Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; Poisoned rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
@@
4 Methodology
To improve poisoning efficiency, it is important to develop proxy-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we will first discuss some key observations.
4.1 Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We contend that the primary reason for the efficiency of backdoor injection lies in the similarity between clean and corresponding poisoned samples. Typically, the labels of poisoned samples are different from those of the original clean samples, meaning that clean and corresponding poisoned samples share similar features but have completely different labels. In this case, samples with high similarity can be viewed as challenging samples in the poisoning task, making them more efficient than low-similarity samples for backdoor attacks. To validate this hypothesis, we conducted empirical studies on the CIFAR-10 dataset using the same experimental settings as in Sec. 3.1. The left part of Figure 4 depicts the similarity distribution of 500 randomly selected samples and 500 FUS-searched samples [39], which demonstrates that FUS-searched efficient data has higher similarity than random samples. We further examined the contribution of similarity to the data efficiency of backdoor attacks by selecting poisoned samples based on similarity. The right part of Figure 4 shows the attack success rate when different poisoned samples were used. As expected, poisoning with high-similarity samples (Top-500 similarity sampling) resulted in a significantly higher attack success rate than poisoning with low-similarity samples (Bottom-500 similarity sampling). Furthermore, we compared random sampling with similarity-selected sampling. However, both poisoning with high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) yielded lower attack success rates than poisoning with random samples (Random sampling). This suggests that similarity can only be used to filter out inefficient samples and has no positive correlation with data efficiency in poisoned attacks. We argue that the degradation of Top-500 similarity samples can be attributed to the overly limited diversity of the poisoned samples. The middle part of Figure 4 shows the t-SNE of the 50K whole dataset and the 500 selected samples, indicating that Top-500 similarity samples have limited diversity. Selecting 500 samples randomly from the 5K samples with the highest similarity (Top-5K similarity random 500 sampling) can effectively mitigate this limitation, as shown in the right part of Figure 4. Among them, Top-5K similarity random 500 sampling outperforms random sampling and achieves competitive results with only a slight cost compared to FUS [39] methods. In summary, the similarity between clean and corresponding poisoned samples and the diversity of poisoned sample set are two key factors in the efficient data of backdoor attacks.
4.2 Threat Model
Attacker’s Capacities. Following the basic requirement of poisoning-based backdoor attacks [22, 39], we assume that attackers are permitted to poison a portion of the training data, but are not privy to any other training components in the victim phase, such as the training loss, schedule, or model structure. These assumptions align with potential real-world scenarios in which such threats can occur.
Attacker’s Goals. Our aim is to enhance the efficiency of poisoning attacks without incurring additional costs, by selecting an appropriate poisoning set from the dataset . This approach is distinct from the current methods that enhance efficiency by designing triggers, and hence, our method can be easily incorporated into other attack technologies at minimal cost.
4.3 Proxy-Free Selection Strategy
@@
Motivated by the observation of forensic features that contribute to efficient data in poisoning-based backdoor attacks, as demonstrated in Section 4.1, we propose a simple yet effective sample selection strategy (PFS) for enhancing poisoning efficiency. Unlike previous studies that rely on a proxy backdoor attack for sample selection, our method is based on the similarity between clean and corresponding poisoned samples, which is only dependent on the Data (D) and Trigger (T), and independent of Data Transform (F) and Hyperparameters (H) (proxy-free). Specifically, we utilize a pre-trained feature extractor to compute the cosine similarity (cos()) in the feature space between clean and corresponding poisoned samples. We then select the top most similar samples and randomly sample from this set as the poisoning set, where is a hyper-parameter that controls the diversity of the poisoned samples. The algorithmic procedure of our method is presented in Algorithm 1.
Moreover, we demonstrate that our method can filter out most of the worthless samples, which also enhances the FUS search [39]. Therefore, we propose a combined approach (FUS+PFS) that utilizes the FUS search on the coarse poisoned set containing samples to further enhance the poisoning efficiency.
5 Experiments
@@
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. 5.2), CIFAR-100 (Sec. 5.3), and Tiny-ImageNet (Sec. A.4 of the Supplementary Materials). We also provide experimental settings in Sec. 5.1, time-consuming analysis in Sec. 5.4, and ablation study in Sec. 5.5.
5.1 Experimental Settings
We conduct experiments with three trigger types (BadNets [13], Blended [4], Optimized [48]), six data transforms (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three DNN architectures (VGG16 [34], ResNet18 [15], PreAct-ResNet18 [16]). We also use four optimizers: SGD-0.01‡‡‡SGD-0.01 means adopting the SGD optimizer with an initial learning rate of 0.01,, SGD-0.02, Adam-0.001, Adam-0.002. The adopted feature extractor is the ResNet-18 that has been pre-trained on the current dataset. Additionally, in Section A.5 of the Supplementary Materials, we present experiments with different pre-trained feature extractors on the CIFAR-10 dataset. Our results show that even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.
We compare our method to the commonly used random selection (Random) [13, 4] and FUS [39]. In the FUS search process, we use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. For our method, we use pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 for all three datasets. Other hyperparameters are kept the same as in Sec. 3.1.
@@ 5.1 Experimental Settings
5.2 Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table 1 presents the attack success rate for BadNets, Blended, and Optimized triggers using different data transforms and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. Our PFS method significantly improves the efficiency of poisoned data for backdoor attacks. In all 54 settings, the attack success rate of the poisoned samples selected by our method is higher than that of randomly selected samples, at the same poisoning ratios. On average, our method outperforms random selection by 0.044, 0.074, and 0.066 for BadNets, Blended, and Optimized triggers, respectively. In general, our method performs better than FUS, achieving stronger attack strength in 48 out of 54 settings. It is worth noting that the FUS search requires hundreds or even thousands of times more time than our method. Combining FUS with our method yields the best results in 51 out of 54 settings, as our method filters out many inefficient samples and narrows the FUS search space.
Results on different optimizer and architectures. Table 2 shows the attack success rate for the BadNets, Blended, and Optimized triggers with different optimizers and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. The results demonstrate the effectiveness of our PFS method. First, for all 36 settings, the attack success rate achieved by our method is higher than that achieved by random selection, with average improvements of 0.036, 0.0437, and 0.017 for the BadNets, Blended, and Optimized triggers, respectively. Second, our method outperforms FUS in 25 out of 36 cases. When combining our method with FUS, we obtain the best results in 33 out of 36 settings.
In the Supplementary Materials, we have included additional experiments to further evaluate the effectiveness and robustness of our proposed strategy. Specifically, in Sec. A.1 and A.2, we present the attack success rate (ASR) and benign accuracy (BA) under poisoned rates of and on the CIFAR-10 dataset. The results demonstrate that our method remains robust to different poisoned rates and target labels while preserving benign accuracy. Additionally, in Sec. A.3, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our findings confirm the effectiveness of our proposed method and support our hypothesis that the efficiency of backdoor attacks is determined by sample similarity and diversity.
@@ 5.2 Experiments on CIFAR-10 Dataset
| PFS (w/o pre-training) | PFS (w pre-training) | FUS () |
| 17 | 562 | 9,150 |
5.3 Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table 3 are similar to those for CIFAR-10. Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
5.4 Analysis of Time Consumption
@@
Table 4 presents a comparison of the running times of our PFS with those of FUS. Our method demonstrates remarkable speed, being faster than FUS. Even with pre-training time included, our method is still approximately faster. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
5.5 Ablation Study
@@
We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoned sample set. The results are shown in Figure 5, and the attack success rates demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the attack success rates also increase. However, too large of a value for can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoned samples, supporting our claim in Sec. 4.1.
6 Conclusion and Future Work
@@
Conclusion. This paper presents empirical evidence and reveals two intriguing phenomena related to data efficiency in backdoor attacks. First, we demonstrate that sample selection methods based on proxy attacks suffer from performance degradation when gaps exist between proxy poisoning attacks and actual victim poisoning processes. Second, we show that the similarity between clean and corresponding poisoned samples, as well as the diversity of the poisoned sample set, are key features for efficient data in backdoor attacks. Based on these findings, we propose a simple yet effective proxy-free sample selection strategy that selects poisoned samples based on the similarity between clean and corresponding poisoned samples. Our method significantly improves the poisoning efficiency on benchmark datasets with almost no additional cost.
@@
Future Work. While the proxy-free sampling strategy we proposed has demonstrated efficient data selection in backdoor attacks, we note that the similarity between clean and corresponding poisoned samples does not necessarily guarantee high efficiency, as similarity only filters out non-efficient samples. In the future, it would be desirable to identify a unified indicator that positively correlates with data efficiency in backdoor attacks.

### v2

@@ IV-B Threat Model
Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; poisoning rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
@@
IV-C Proxy Attack-free Selection Strategy
@@
Drawing inspiration from our analysis of forensic features contributing to efficient data in poisoning-based backdoor attacks, as showcased in Sec. IV-A, we introduce a simple yet efficient sample selection strategy termed the Proxy attack-Free Strategy (PFS) for improving poisoning efficiency. Unlike prior approaches that rely on a proxy attack task for sample selection, our technique capitalizes on the similarity between benign and corresponding poisoning samples, which is only related to the Data (D) and Trigger (T), liberating it from dependence on the settings of actual attack within victim phase. Specifically, we adopt a pre-trained feature extractor to compute the cosine similarity (cos()) within the feature space between benign and corresponding poisoning samples. Subsequently, we identify the top most similar samples and randomly select from this subset as the poisoning set, where is a hyper-parameter that modulates the diversity of the poisoning samples. The procedure of our methodology is presented in Algorithm 1. Acknowledging that our PFS relies on a feature extractor pre-trained on a clean dataset to compute similarity, we contend that the essence of FUS and PFS diverges significantly. Our PFS identify the forensic features of efficient samples in poisoning-based backdoor attacks, and adopt the similarity and diversity to identify efficient samples, which is disentangled from the actual attack within victim phase. Therefore, the effectiveness of PFS primarily depends on the accuracy of similarity measurement.
Furthermore, we confirm that our method efficiently filters out most of the ineffective samples, thus enhancing the performance of FUS [50]. Therefore, we introduce a collaborative strategy called FUS+PFS, which combines FUS search with the initial coarse poison set containing samples to further enhance poisoning efficiency.
IV-D Theoretical Analyses
Additionally, we provide theoretical analyses of the proposed PFS in terms of both active learning and neural tangent kernel.
Perspective of Active Learning. The process of selecting efficient poisoning samples in the context of backdoor attacks can be analogized to the principles of active learning. Active learning (AL) is a machine learning paradigm that efficiently acquires annotated data from an extensive pool of unlabeled data [35]. Its objective is to prioritize human labeling efforts on the most informative data points, leading to improved model performance while reducing data annotation costs. Similarly, the task of selecting efficient poisoning samples in a backdoor attack aims to identify the most informative data points that can optimize the efficiency of the poisoning process, thereby minimizing the costs for the attacker. A pivotal aspect of active learning encompasses the concepts of uncertainty and diversity in the dataset [6]. Methods that rely on uncertainty utilize the predictive confidence of model to select challenging examples, while diversity-based sampling leverages the dissimilarity among data points, often involving clustering techniques. Notably, the combination of uncertainty and diversity strategies has demonstrated significant success in active learning [56, 42]. Similarly, our proposed PFS takes into account both uncertainty (Line 8 in Algorithm 1) and diversity (Line 9 in Algorithm 1) of data points.
Furthering our understanding of data point uncertainty within backdoor attacks, we delve into more comprehensive analyses to underscore the connection between the proposed measure of similarity between benign and corresponding poisoning samples and the expression of data point uncertainty. Drawing from theoretical insights provided by margin theory in active learning [2], we recognize that examples near the decision boundary offer significant potential for reducing annotation requirements. In the context of a backdoor attack, we exclusively consider the true class and the attack-target class of the clean sample and its corresponding poisoning sample , respectively, forming a two-class classification task. For any given poisoning sample , its corresponding clean sample is the nearest sample, albeit labeled with a different category. This configuration establishes a margin , with the decision boundary residing centrally within the margin. As such, the distance between the sample and the decision boundary is proportionate to . To summarize, the distance between benign and corresponding poisoning samples, , can be employed to convey the uncertainty of data points. This assessment is carried out using a pre-trained feature extractor to compute distances within the feature space. Figure 6 illustrates the effective integration of individual similarity and set diversity in constructing a poisoning injection.
Perspective of Neural Tangent Kernel (NTK). To further explain this interesting phenomenon (i.e., similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the efficiency of data in backdoor attacks), we exploit recent studies on NTK [22, 15, 16, 30] for analyzing the sample efficiency of backdoor attacks:
Theorem 1.
Suppose the training dataset consists of benign samples and poisoned samples , whose images are i.i.d. sampled from uniform distribution and belonging to classes. Assume that the DNN is a multivariate kernel regression and is trained via . For the expected predictive confidences over the target label , we have: , where is poisoning testing samples of attacks.
In general, Theorem 1 suggests that selecting samples characterized by high similarity between clean () and corresponding poisoning samples (), as well as high similarity between poisoning samples () and poisoned testing samples (), enhances the confidence in predicting poisoned samples to the target class. In the context of poisoning-based backdoor attacks, where training samples share the same distribution as testing samples, high similarity between poisoning samples and poisoned testing samples indicates congruence in the distribution between the poisoning set and the training set, signifying a high level of diversity within the poisoning set. The proof of this theorem can be found in the Sec. A of the Supplementary Materials.
V Experiments
@@
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. V-B), Tiny-ImageNet (Sec. V-C), and CIFAR-100 (Sec. E of the Supplementary Materials). We also provide experimental settings in Sec. V-A, comparison with recent sample selection methods in Sec. V-D, time consumption analysis in Sec. V-F, ablation study in Sec. V-G, and experiment against defense method in Sec. V-H.
@@
V-A Experimental Settings
We conduct comprehensive experiments encompassing three trigger types§§§Additionally, we consider the recent backdoor attacks such as WaNet [38] and ISSBA [27] in Sec. V-D. (BadNets [14], Blended [4], Optimized [61]), six data transformations (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three different DNN architectures (VGG16 [43], ResNet18 [19], PreAct-ResNet18 [20]). Furthermore, we explore the effects of four optimizers: SGD-0.01¶¶¶SGD-0.01 implies the adoption of the SGD optimizer with an initial learning rate of 0.01., SGD-0.02, Adam-0.001, and Adam-0.002. The employed feature extractor is ResNet18, pre-trained on the specific dataset using the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, along with two standard data augmentations: random crop and random horizontal flip. Each training iteration spans 70 epochs, conducted with a batch size of 256. In addition, Sec. V-E presents experiments involving various pre-trained feature extractors on the CIFAR-10 dataset. Our findings reveal that even utilizing a commonly used ImageNet pre-trained feature extractor yields notable enhancements compared to the baseline.
We perform a comprehensive comparative analysis, evaluating our method against the widely employed random selection (Random) [14, 4] and FUS [50] approaches. In the FUS search procedure, we utilize the VGG16 architecture, coupled with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer to establish the proxy attack. Conversely, for our proposed method, we utilize pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 across all three datasets. All other hyperparameters remain consistent with those outlined in Sec. III-A. Additionally, Sec. V-D shows the comparison with recent sample selection methods (FUS [50], OPS [17], Gradient[10], RD Score[49]) on the Tiny-ImageNet dataset.
V-B Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table I provides a comprehensive overview of the attack success rates for BadNets, Blended, and Optimized triggers, considering various data transforms and architectures. Constructing the proxy attack for FUS, we employ VGG16 with RandomCrop+RandomHorizontalFlip and SGD-0.01 as the optimizer. Our proposed PFS method significantly enhances the poisoning efficacy in backdoor attacks. Across all 54 configurations, the attack success rate of poisoning samples selected through our approach consistently surpasses that of randomly selected samples at identical poisoning rates. On average, our method achieves a improvement of 0.044, 0.074, and 0.066 in terms of attack success rate for BadNets, Blended, and Optimized triggers, respectively. In general, our method outperforms FUS, demonstrating greater ASR in 48 of the 54 scenarios. Notably, FUS searches necessitate significantly more time compared to our method, often by hundreds or even thousands of times. Furthermore, when the model is VGG16 and the trigger is set to Blended, FUS surpasses our PFS. This can be attributed to two factors: (i) Both the proxy attack model employed in FUS and the model utilized in the victim phase are VGG16; (ii) In comparison to BadNets and Optimized triggers, the Blended trigger increases the similarity distance between benign and corresponding poisoning samples, thereby marginally impacting the performance of our proposed PFS. The combination of FUS and PFS, denoted as FUS+PFS, excels in 51 out of the 54 scenarios. This outcome stems from PFS effectively eliminating inefficient samples, consequently decreasing the search space for FUS.
@@ V-B Experiments on CIFAR-10 Dataset
Results on different optimizer and architectures. Table II presents the ASR for the BadNets, Blended, and Optimized triggers, considering different optimizers and architectures. We utilize VGG16 with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer for constructing the proxy attack in FUS. The outcomes underscore the effectiveness of our PFS method. Firstly, across all 36 configurations, our PFS consistently achieves higher attack success rates than random selection, with average improvements of 0.0437, 0.052 and 0.017 for the BadNets, Blended, and Optimized triggers, respectively. Secondly, our PFS outperforms FUS in 25 out of the 36 cases. Similar to previous experiments, when the model is VGG16 and the trigger is set to Blended, FUS outperforms our PFS. Notably, the combination of our method with FUS, referred to as FUS+PFS, achieves the best performance in 33 out of the 36 settings.
The Supplementary Materials contain additional experiments aimed at further evaluating the effectiveness and robustness of our proposed strategy. Specifically, in Sec. B and Sec. C, we provide the attack success rate (ASR) and benign accuracy (BA) under poisoning rates of and on the CIFAR-10 dataset, respectively. The outcomes illustrate the robustness of our method across different poisoning rates while maintaining benign accuracy. Moreover, in Sec. D, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our results consistently affirm the effectiveness of the proposed pfs and support our hypothesis that the efficiency of backdoor attacks depends on the interaction between individual similarity and set diversity.
V-C Experiments on Tiny-ImageNet Dataset
Results on different data transforms and architectures. Table III shows that PFS outperforms the random method in terms of attack success rates of poisoning samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. In contrast to experiments on the CIFAR-10 dataset, our proposed PFS method consistently achieves higher average ASR than FUS across different data transformations on the Tiny-ImageNet dataset. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method on the scenarios involving extensive datasets.
Results on different data transforms, optimizer, and architectures. To further illustrate the effectiveness of our proposed PFS, Figure 7 visually shows the average ASR across various data transforms and optimizers on the Tiny-ImageNet dataset. The findings indicate that all advanced sample selection strategies (FUS, PFS, and FUS+PFS) achieve notably superior results compared to the random selection strategy (RSS). Furthermore, our proposed PFS consistently achieves better ASR compared to FUS across all settings.
@@ V-C Experiments on Tiny-ImageNet Dataset
| Dataset | PFS (w/o pre-training) | PFS (w pre-training) | FUS |
| CIFAR-10 | 17 | 562 | 9,150 |
| Tiny-ImageNet | 19 | 6784 | 106284 |
V-D Comparison with Recent Sample Selection Strategies
Table IV presents a comparative analysis between our PFS and recent sample selection strategies (FUS [50], OPS [17], Gradient[10], and RD Score[49]) on the Tiny-ImageNet dataset. Various recent triggers (Blended, WaNet, and ISSBA) are considered in this experiment. The results demonstrate that all sample selection strategies exhibit superior ASR compared to the random selection strategy (RSS), with our proposed PFS consistently demonstrating the highest sample efficiency across all settings when compared to other recent sample selection strategies.
V-E Experiments on Different Pre-trained Feature Extractor
Although both FUS and PFS rely on a proxy task, there are essential differences between proxy tasks. FUS relies on a proxy attack task closely resembling the actual attack. Nevertheless, due to the inherent shortcuts of backdoor learning, overfitting to the data transformation and hyperparameter of proxy attack significantly impact actual injection efficiency. In contrast, our PFS selects efficient samples considering individual similarity and set diversity. As a result, PFS operates independently of the specific attack deployed during the victim phase, with minimal impact from data transformation and hyperparameters within the pre-trained feature extractor on sample efficiency during this phase. This independence is evidenced in our experiments with various pre-trained extractors, as illustrated in Figure 8. Remarkably, even a common extractor pre-trained on ImageNet results in considerable improvements compared to the baseline.
V-F Analysis of Time Consumption
@@
Table V presents a comparison of the running times of our PFS with those of FUS. PFS demonstrates remarkable speed, being faster than FUS on CIFAR-10 dataset. Even with pre-training time included, our method is still approximately faster on CIFAR-10 dataset. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
V-G Ablation Study
@@
We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoning sample set. The results are shown in Figure 9, and the attack success rates demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the attack success rates also increase. However, too large of a value for can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoning samples, supporting our claim in Sec. IV-A. In our PFS, we set in all experimental settings.
V-H Experiment with Backdoor Defense.
We evaluate the robustness of PFS against backdoor defense using the pruning-based defense method [32] and the tune-based defense method FST [36]. Pruning [32] is a widely used defense mechanism. It achieves this by pruning neurons that remain inactive in response to benign inputs, thereby neutralizing the effectiveness of the backdoor behavior. On the other hand, FST [36] is a tuning-based backdoor purification method. It operates by inducing feature shifts through deliberate adjustments to the classifier weights, effectively diverging them from the initially compromised weights. As depicted in Figure 10, PFS demonstrates superior robustness against Pruning [32] compared to RSS and FUS. Similarly, Figure 11 illustrates that our proposed PFS effectively against the tune-based defense method FST [36] compared to RSS and FUS. It is noteworthy that while FST [36] demonstrates superior defense capabilities compared to Pruning [32], it comes at the expense of a substantial reduction in benign accuracy.
VI Conclusion and Future Work
@@
Conclusion. This paper contributes empirical insights into two interesting phenomena associated with data efficiency in backdoor attacks. Firstly, we highlight the performance degeneration experienced by sample selection methods relying on proxy attacks when discrepancies arise between proxy attacks within attacker phase and actual poisoning processes within the victim phase. Furthermore, we argue that the significance of both the similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set as essential factors for efficient data in backdoor attacks. Leveraging these observations, we introduce a simple yet efficient proxy attack-free sample selection strategy that is driven by the similarity between benign and corresponding poisoning samples. Our method significantly improves poisoning efficiency across benchmark datasets with minimal additional cost.
@@
Future Work. While the proposed PFS has demonstrated efficient data selection in backdoor attacks, we note that similarity between benign and corresponding poisoned samples does not necessarily guarantee efficiency, as similarity only filters out non-effective samples. In the future, it would be desirable to identify an unified indicator that positively correlates with data efficiency in backdoor attacks. Moreover, our proposed PFS is customized specifically for dirty-label backdoor attacks. In future research, we intend to explore efficient sample selection strategies designed specifically for clean-label backdoor attacks.
Acknowledgments
The work was supported in part by the National Natural Science Foundation of China under Grands U19B2044 and 61836011.